1,305 views

2 Answers

Best answer
4 4 votes

Uniform random variable X with parameter a,b:

F(X)=$\frac{1}{b-a}$

F(X)=2 ,a=.5 and b= 1.

To find probability distribution of the interval (0.55,0.60) and (0.75 , 0.85), find the area of the region they cover.

Probability (X)= area 1 + area 2

area 1 :- As it will make rectangle with height=2 and width=0.60-0.55=0.05. Area1= 0.05*2=0.1

area 2 :- 

As it will make rectangle with height=2 and width=0.85-0.75=0.1. Area1= 0.1*2=0.2

Probability = 0.1+0.2=0.3

https://en.wikipedia.org/wiki/Uniform_distribution_(continuous)

• selected by
0 0 votes
1.00-0.50= 0.50 (Total case)

0.60 to 0.55 = 0.05, 0.85-0.75=0.10

0.05+0.10=0.15 (favorable case)

Probability = $\frac{0.15}{0.50}$=0.3
Position:
Show:

Related questions

0 0 votes
1 1 answer
471
471 views
32 32 votes
3 answers 3 answers
15.7k
15.7k views
gatecse asked Sep 21, 2014
15,650 views
Let $f(x)$ be the continuous probability density function of a random variable $x$, the probability that $a < x \leq b$, is :$f(b-a)$$f(b) - f(a)$$\int\limits_a^b f(x) dx...
7 7 votes
2 2 answers
544
544 views
GO Classes asked May 5
544 views
Let $X$ be an exponentially distributed random variable with mean $1/\ln 3$. Define $Y=\lfloor X \rfloor$, where $\lfloor X \rfloor$ is the greatest integer less than or ...
2 2 votes
1 1 answer
279
279 views
GO Classes asked May 1
279 views
Let $X_1,X_2,\ldots,X_n$ be independent discrete uniform random variables on ${1,2,\ldots,n}$. That is, $P(X_i=j)=\frac{1}{n}$ for each $i$ and each $j\in{1,2,\ldots,n}$....